OpenStudy (anonymous):

A sphere just fits inside a cube with edges 18cm long. What is the volume of the sphere?

Directrix (directrix):

If the sphere is an exact fit for the cube, then the faces of the cube are tangent to the sphere. The diameter of the sphere would be 18 and its radius 9. The volume of a sphere is given by V = (4/3) * pi* r^3 V = (4/3) * pi * 9 ^ 3 V= (4/3) * pi * 729 Multiply those numbers to get an answer in cubic centimeters. @queenaliya